[10] 1. Solve the first-order equation Tư + (x – 2) = -xe-3, > 0.
Answer #1

Queation Page-t solve the follo ollowing firet order we have to equation, ayt (x-2) y se. et ; ayo (1-2) y = - e dy dae + Papage-2 or g. 10 -S da &K yine se? e da + K -S ye. ont a + k 4 oC 018 = set +K 4 06 28 to -4 t.tk y= K + some other constant.